# GaussianDiffusion: 3D Gaussian Splatting for Denoising Diffusion Probabilistic Models with Structured Noise

### v2

@@ Abstract
Text-to-3D, known for its efficient generation methods and expansive creative potential, has garnered significant attention in the AIGC domain. However, the pixel-wise rendering of NeRF and its ray marching light sampling constrain the rendering speed, impacting its utility in downstream industrial applications. Gaussian splatting has recently shown a trend of replacing the traditional pointwise sampling technique commonly used in NeRF-based methodologies, and it is changing various aspects of 3D reconstruction. This paper introduces a novel text to 3D content generation framework based on Gaussian splatting and produces more realistic renderings. The challenge of achieving multi-view consistency in 3D generation significantly impedes modeling complexity and accuracy. Taking inspiration from SJC, we explore employing multi-view noise distributions to perturb images generated by 3D Gaussian splatting, aiming to rectify inconsistencies in multi-view geometry. We ingeniously devise an efficient method to generate noise that produces Gaussian noise from diverse viewpoints, all originating from a shared noise source. Furthermore, vanilla 3D Gaussian-based generation tends to trap models in local minima, causing artifacts like floaters, burrs, or proliferative elements. To mitigate these issues, we propose the variational Gaussian splatting technique to enhance the quality and stability of 3D appearance. To our knowledge, our approach represents the first comprehensive utilization of Gaussian splatting across the entire spectrum of 3D content generation processes.
@@ 1 Introduction
  • We introduce structured noise from various viewpoints for the first time, aiming to tackle the challenge of maintaining multi-view geometric consistency, such as the problem of multi-faceted structures, through a noise injection approach. Furthermore, we propose an effective method for generating noise based on Gaussian Splatting.
@@ 3 Method
So, for a given pixel position, C can be represented as a standard normal distribution. That is $P(C|U,\pi)=N(0,\text{Var}(C))$ can be standardized to $N(0,I)$, allowing for noise to be added to the novel view synthesized by Gaussian Splatting. This noise is associated with the viewpoint $\pi$ and does not affect the numerical stability of $N(0,I)$ noises added by SJC, as shown in Figure  3

### v3

@@ Abstract
Text-to-3D, known for its efficient generation methods and expansive creative potential, has garnered significant attention in the AIGC domain. However, the pixel-wise rendering of NeRF and its ray marching light sampling constrain the rendering speed, impacting its utility in downstream industrial applications. Gaussian Splatting has recently shown a trend of replacing the traditional pointwise sampling technique commonly used in NeRF-based methodologies, and it is changing various aspects of 3D reconstruction. This paper introduces a novel text to 3D content generation framework, Gaussian Diffusion, based on Gaussian Splatting and produces more realistic renderings. The challenge of achieving multi-view consistency in 3D generation significantly impedes modeling complexity and accuracy. Taking inspiration from SJC, we explore employing multi-view noise distributions to perturb images generated by 3D Gaussian Splatting, aiming to rectify inconsistencies in multi-view geometry. We ingeniously devise an efficient method to generate noise that produces Gaussian noise from diverse viewpoints, all originating from a shared noise source. Furthermore, vanilla 3D Gaussian-based generation tends to trap models in local minima, causing artifacts like floaters, burrs, or proliferative elements. To mitigate these issues, we propose the variational Gaussian Splatting technique to enhance the quality and stability of 3D appearance. To our knowledge, our approach represents the first comprehensive utilization of Gaussian Diffusion across the entire spectrum of 3D content generation processes.
@@ 1 Introduction
  • We introduce structured noise from various viewpoints for the first time, aiming to tackle the challenge of maintaining multi-view geometric consistency, such as the problem of multi-faceted structures, through a structured Gaussian noise injection approach. Furthermore, we propose an effective method for generating structured noise based on the Gaussian diffusion process.
@@ 3 Method
So, for a given pixel position, C can be represented as a standard normal distribution. That is $P(C|U,\pi)=N(0,\text{Var}(C))$ can be standardized to $N(0,I)$, allowing for noise to be added to the novel view synthesized by Gaussian Splatting. This noise is then used in the subsequent diffusion noise-adding and denoising process, the Gaussian diffusion process. This noise is associated with the viewpoint $\pi$ and does not affect the numerical stability of $N(0,I)$ noises added by SJC, as shown in Figure  3